# SurgicAI: A Fine-grained Platform for Data Collection and Benchmarking in Surgical Policy Learning

### v1

@@ A.2 Image-Based Imitation Learning Settings
Training utilizes a dataset that includes visual and proprioceptive data paired with corresponding actions. The high-level policy is updated using cross-entropy loss to generate discrete classifications, while the low-level policy aims to minimize the mean squared error between the expert action and the output action. Both levels use the ADAM optimizer with a learning rate of 0.001.
A.3 TD3+HER+BC
One SOTA RL technique used for LLPs is TD3 combined with HER, BC, and Q-filter as shown in Fig. 4. HER allows the agent to learn from unsuccessful attempts by reinterpreting these attempts as successful ones for different goals, thereby generating additional useful experiences. Specifically, for each real transition, three virtual transitions are created by sampling new goals, significantly increasing the amount of meaningful training data.
Figure 4: Workflow of the TD3+HER+BC for policy learning. The rollout collected from SRC environment is stored in the form $\left(s_{t},a_{t},s_{t+1},r,ag,dg\right)$. Both HER module and expert demonstrations enrich the buffer with additional transitions. The TD3 network, comprising actors and critics, updates its policy based on the rollout from expanded buffer. The behavior cloning loss $\left(L_{BC}\right)$ minimizes the error between predicted and expert actions, while using a $\mathrm{Q}$ filter to ensure the policy remains close to the expert while avoiding suboptimal solutions.
@@ A.3 TD3+HER+BC
A.4 Weight Settings of BC Loss
A.4.1 Calculation of RL Loss
@@
$L(\theta)=\frac{1}{N_{1}}\sum_{i=1}^{N_{1}}\left[\left(Q_{\theta}(s_{i},a_{i})-y_{i}\right)^{2}\right]$
@@
$y_{i}=r_{i}+\gamma\min_{j=1,2}Q_{\theta_{j}^{\prime}}(s_{i}^{\prime},\pi_{\phi^{\prime}}(s_{i}^{\prime}))$
Here, $\gamma$ is the discount factor, $Q_{\theta_{j}^{\prime}}$ are the target critic networks, $\pi_{\phi^{\prime}}$ is the target actor network, and the expectation is approximated by samples from the replay buffer.
@@
$\nabla_{\theta_{\pi}}J(\pi)=\frac{1}{N_{1}}\sum_{i=1}^{N_{1}}\nabla_{a}Q(s,a|\theta_{Q})\bigg{|}_{s=s_{i},a=\pi(s_{i})}\nabla_{\theta_{\pi}}\pi(s|\theta_{\pi})\bigg{|}_{s_{i}}$
A.4.2 Calculation of BC Loss
@@
A.4.3 Integration of Losses with Weighted Sum
To effectively combine the RL and BC losses, a weighted combination [28] is used to update actor policy, where $\lambda_{1}$ and $\lambda_{2}$ adjust the influence of the RL and BC losses respectively:
@@ A.4.3 Integration of Losses with Weighted Sum
A.5 Heuristic Trajectory collection
A.5.1 Data Collection and Preprocessing
Data Collection
To obtain expert data, we collect the demonstrations from human subjects. We ask the human subjects to perform a single-throw suture procedure in the simulation environment[37] using dVRK MTM. We record all the kinematic and dynamic data during the teleoperation. Then, taking advantage of the great repeatability of AMBF[38], we can steadily replay the movements and select specific kinds of data we would like to collect. The whole data collection framework follows the pipeline developed in[24] :
Data Decouple
@@ A.5.1 Data Collection and Preprocessing
Data Filter
To obtain solid motion data, it is also important to exclude the static states to reduce the variance. Therefore, after data decouple, we also implement a velocity-based filter to the segmented sub-dataset. We set the thresholds of 0.2 mm/s and 0.1 deg/s for linear and angular velocity lower limits. For any data points which have velocities smaller than the lower limits, the data points will be considered to be static and removed from the dataset.
A.5.2 Imitation-learning based solution
@@
The whole pipeline acheiving the suturing task automation in Zhou et. al. can be summarized in the following pseudo algorithm:
  1.Data Collection, data decouple and data preprocessing.
  2.Select one needle trajectory from expert demonstrations, calculated the optimal learning weights for DMP using locally weighted regression[57].
@@
  5. Convert the needle trajectory into the PSM trajectory using the inverse kinematic.

### v3

@@ A.2 Image-Based Imitation Learning Settings
The training utilizes a dataset that includes visual and proprioceptive data paired with corresponding actions. The high-level policy is updated using cross-entropy loss to generate discrete classifications, while the low-level policy aims to minimize the mean squared error between the expert action and the output action. Both levels use the ADAM optimizer with a learning rate of 0.001.
A.3 Wall Time Report
We evaluated the average frames per second (FPS) and elapsed time for TD3+HER+BC after reaching 150,000 total time steps in our soft body simulation. The results showed an FPS of approximately 120 Hz, with an elapsed time of around 1,696 seconds. These experiments were conducted in headless mode using a Tesla T4 GPU on the Microsoft Azure platform.
To further accelerate training, we are investigating methods such as parallelizing SRC simulations and sampling data concurrently. While we are still refining these techniques to optimize sampling efficiency, we have provided prototype demonstrations in the repository. This notebook allows users to initiate and manage independent environments in parallel. Details of these developments will be included in future revisions, and we are actively working on integrating SRC with Isaac Sim [66] to enhance overall efficiency.
A.4 TD3+HER+BC
One SOTA RL technique used for LLPs is TD3 combined with HER, BC, and Q-filter as shown in Figure 4. HER allows the agent to learn from unsuccessful attempts by reinterpreting these attempts as successful ones for different goals, thereby generating additional useful experiences. Specifically, for each real transition, three virtual transitions are created by sampling new goals, significantly increasing the amount of meaningful training data.
Figure 4: Workflow of the TD3+HER+BC for policy learning. The rollout collected from the SRC environment is stored in the form $\left(s_{t},a_{t},s_{t+1},r,ag,dg\right)$. Both HER module and expert demonstrations enrich the buffer with additional transitions. The TD3 network, comprising actors and critics, updates its policy based on the rollout from the expanded buffer. The behavior cloning loss $\left(L_{BC}\right)$ minimizes the error between predicted and expert actions, while using a $\mathrm{Q}$ filter to ensure the policy remains close to the expert while avoiding suboptimal solutions.
@@ A.4 TD3+HER+BC
A.5 Weight Settings of BC Loss
A.5.1 Calculation of RL Loss
@@
$L(\phi)=\frac{1}{N_{1}}\sum_{i=1}^{N_{1}}\left[\left(Q_{\phi}(s_{i},a_{i})-y_{i}\right)^{2}\right]$
@@
$y_{i}=r_{i}+\gamma\min_{j=1,2}Q_{\phi_{j}^{\prime}}(s_{i}^{\prime},\pi_{\phi^{\prime}}(s_{i}^{\prime}))$
Here, $\gamma$ is the discount factor, $Q_{\phi_{j}^{\prime}}$ are the target critic networks, $\pi_{\phi^{\prime}}$ is the target actor network, and the expectation is approximated by samples from the replay buffer.
@@
$\nabla_{\phi_{\pi}}J(\pi)=\frac{1}{N_{1}}\sum_{i=1}^{N_{1}}\nabla_{a}Q(s,a|\phi_{Q})\bigg{|}_{s=s_{i},a=\pi(s_{i})}\nabla_{\phi_{\pi}}\pi(s|\phi_{\pi})\bigg{|}_{s_{i}}$
A.5.2 Calculation of BC Loss
@@
A.5.3 Integration of Losses with Weighted Sum
To effectively combine the RL and BC losses, a weighted combination [31] is used to update the actor policy, where $\lambda_{1}$ and $\lambda_{2}$ adjust the influence of the RL and BC losses respectively:
@@ A.5.3 Integration of Losses with Weighted Sum
A.6 Heuristic Trajectory collection
A.6.1 Data Collection and Preprocessing
Data Collection To obtain expert data, we collect the demonstrations from human subjects. We ask the human subjects to perform a single-throw suture procedure in the simulation environment [18] using a dVRK MTM. We record all the kinematic and dynamic data during the teleoperation. Then, taking advantage of the great repeatability of AMBF [41], we can steadily replay the movements and select specific kinds of data we would like to collect. The whole data collection framework follows the pipeline developed in [27] as shown in Figure 5.
Figure 5: Expert demonstration data collection pipeline.
Data Decoupling
@@ A.6.1 Data Collection and Preprocessing
Data Filter To obtain solid motion data, it is also important to exclude the static states to reduce the variance. Therefore, after data decoupling, we also implement a velocity-based filter to the segmented sub-dataset. We set the thresholds of 0.2 mm/s and 0.1 deg/s for linear and angular velocity lower limits. For any data points which have velocities smaller than the lower limits, the data points will be considered to be static and removed from the dataset.
A.6.2 Imitation-learning based solution
@@
The whole pipeline achieving the suturing task automation in Zhou et. al. can be summarized in the following algorithm:
  1.Data Collection, data decoupling and data preprocessing.
  2.Select one needle trajectory from expert demonstrations, calculate the optimal learning weights for DMP using locally weighted regression[71].
@@
  5. Convert the needle trajectory into the PSM trajectory using the inverse kinematics.